SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (June 14, 2005)

Operations

  • June 13 (Monday): L1 trigger rate went up to 20kHz and >70% dead time, mostly from EMC trigger. ODF group got paged first perhaps because they were working on the master crate earlier today. Jamie looked at the trg fastmon plots which showed that some of the output lines were dropped (eg. in run 55259) and it looked like a glt configuration problem. He then correlated that to the createFileTable from the release based on offline 18.0.4 which didn't have the TrgParser fix in yet, but was installed Monday morning. Rainer pointed out that the fix is in 18.0.4d, 18.1.0 and later. Jim put the createFileTable link back to the previous working version and problem was resolved. Runs 55259-55266 are bad.

EMT

  • Dan is working on understanding how to put the prescale factor in for EMT in the configuration database. This is one way to reduce dead time from EMT. Walt pointed out in last week's detector meeting that the EMT contibutes about ~0.2% to dead time, and that we lost about 20 min/week of data due to this.
  • We discussed Dan's outline of the code for the efficiency plots which will be in L1TFMonEmt.

DCT/DCZ

  • The TSF board that Ming sent out arrived UK this week.
  • Jamie posted a HN about a potential problem with the ZPD distribution. .ps and .ppt. We had a discussion of this during this meeting. Su Dong suspected that it was related to the spikes in SL2. Kris mentioned that he didn't see anything wrong in the efficiency from the OPR plots.

Data Quality and Monitoring

  • The fastmon plots for trigger crate/board damage check have been in JAS and .ps archive since last Friday. There are still some debugging to do because the TsfStatusDigit didn't get put in the event in the version of fastmon app that runs at IR2.
  • The BLT/ZPD event struction plots are also in OEP for JAS fastmon since last Friday. Note that previously, it was the BLT/PTD was in there and now it was taken out and replaced by the BLT/ZPD.
  • Kris updated the reference L1T fastmon JAS plots. He noticed that the plots for the ZPD rho occupancy looked different in live fastmon and offline fastmon running on L1TFMonXtcApp. Su Dong pointed out that the online FastMon looks at all L1 triggers while previously the reference plots were rerun against stored .xtc file which is against events passing L3. Therefore, shifters have always been seeing a difference between the data and reference plots, especially for the ZPD plots such as the SL7 and SL10 plots which we previously thought was due to different beam background condition. Kris copied the histrograms from live fastmon to make a set of reference plots to use in JAS, and the data and reference plots now agree a lot better.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.